OpenStudy (help2001):

MEDAL AND FAN!!!!!!!!!11 Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 five over 3 comma 4 over 3 negative 5 over 3 comma negative 4 over 3 negative 3 over 5 comma negative 3 over 4 3 over 4 comma 3 over 5

OpenStudy (help2001):

@igigighjkl @Brill

OpenStudy (anonymous):

Ok so let's do this shall we So in the second equation it says y=2x+2 right?

OpenStudy (help2001):

yes

OpenStudy (help2001):

now what? @Brill

OpenStudy (help2001):

hurry a little i have to wake up at 4 am lol and then on the same day stay up till 3 am

OpenStudy (anonymous):

Right so then you put it into the other one and that means\[x+(2x+2)=3\]Solve for x please

OpenStudy (help2001):

ok so z is -3 right?

OpenStudy (help2001):

i meant x @Brill

OpenStudy (help2001):

ok

OpenStudy (help2001):

hello 2brill?

OpenStudy (therawbugeyes):

hey guys

OpenStudy (anonymous):

Ok so let's subtract 2 from both sides 3x=-5 x=-5/3 Let's sub -5/3 for x in either equation -5/3+y=-3 y=-4/3